OpenStudy (anonymous):

the legend of a map says that .5 inches is equal to 10 miles on a map. if a highway on a map is 1.75 inches long, how long is the actual hughway

OpenStudy (anonymous):

Do a ratio.

zepdrix (zepdrix):

Since we're using the same map for both measurements, the ratio of inches to miles will be the same for both quantities. So we can set it up as a ratio and compare. \[\frac{ .5inches }{ 10miles }=\frac{ 1.75inches }{ x miles }\]

OpenStudy (anonymous):

.5/10 = 1.75/X